OpenStudy (anonymous):

how is 4/3 hours equal to 1hr20min?????

OpenStudy (anonymous):

60 minutes in an hour. A third of 60 is 20, so 4/3 of 60 is 80, which is an hour and 20 minutes.

OpenStudy (anonymous):

i second that :D

OpenStudy (anonymous):

happy???????

OpenStudy (anonymous):

who?

OpenStudy (anonymous):

3/4 is not as same as 4/3.right???

OpenStudy (anonymous):

in question i asked i get the answer as this 1/2 + 1/4 = 3/4 and it says 3/4 hrs= 1hr 20min so i was confused can tell me the answer

OpenStudy (anonymous):

does 3/4 and 4/3 are same??

OpenStudy (anonymous):

no right

OpenStudy (anonymous):

check the website and the last answer in that

OpenStudy (anonymous):

@dwshean

OpenStudy (anonymous):

all question are to u sir

OpenStudy (anonymous):

hellllooooooooooooooo anybody here?

OpenStudy (anonymous):

You have to think of it in terms of time. His answer is right saying that in 1 hour, you completed 3/4 of a whole truck. However, the question asks how much time it takes to complete the whole truck. So they still need to do 1/4 of a truck. Since they did 3 of the 4 parts per 60 minutes, the 1 part they have left will be 60/3, which is 20. So add up your time and that is 1 hour and 20 minutes, or 4/3 hours.